OpenStudy (anonymous):

The sum of the squares of three negative numbers p,q,r equals the sum of the products of all possible pairs of the three numbers. Which of the following equals p3qr2 ? 1) (pq + qr + rp)3 2) q2(p3r + rp3)/2

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

we're given : \(p^2+q^2+r^2 = pq=qr=rp\) so, \(p^3qr^2 = pq*(pr)^2 = (p^2+q^2+r^2)^3\)

OpenStudy (anonymous):

Right answer id option 2 @ganeshie8

OpenStudy (turingtest):

"The sum of the squares of three negative numbers p,q,r equals the *sum* of the products of all possible pairs of the three numbers." I think we are given that\[p^2+q^2+r^2=pq+qr+rp\]I could be wrong; I'm a bit rusty...

hartnn (hartnn):

yes, ^^

OpenStudy (anonymous):

Which of the following equals p3qr2 ?@TuringTest

OpenStudy (turingtest):

I'm pretty bad at these problems, but\[p^2=pq+qr+rp-q^2-r^2\]\[r^2=pq+qr+rp-q^2-p^2\]\[pq=p^2+q^2+r^2-qr-rp\]multiplying all that out should give an answer, but I'm sure there's a smarter way to do it.

OpenStudy (anonymous):

Thank you :)

OpenStudy (turingtest):

You're welcome

hartnn (hartnn):

something is off.. the answer is q2(p3r + rp3)/2 <<p^3 r and r p^3 are same!!>> which means p^3 q r^2 = q^2(p^3 r + rp^3)/2 holds true p^3 and q can be factored out from the right and gets cancelled, r^2 = q (r+r)/2 = rq r= q holds true ...